OpenStudy (kabase1234):

Please help! I'll fan and medal Find an equation for the nth term of the sequence. -3, -12, -48, -192, ... an = 4 • -3n + 1 an = -3 • 4n - 1 an = -3 • 4n an = 4 • -3n

OpenStudy (kabase1234):

@freckles

OpenStudy (kabase1234):

@mathmate @yanasidlinskiy

OpenStudy (kabase1234):

Please help

OpenStudy (shamim):

.=multiplication?

OpenStudy (freckles):

ask yourself these questions: what pattern do you notice? is this a geometric or arithmetic sequence? is there a common ratio or a common difference? what can i do to get from a previous term to the term after?

OpenStudy (kabase1234):

Can you help me solve it?

OpenStudy (freckles):

I think you have not written your choices correctly

OpenStudy (kabase1234):

That's how they were

OpenStudy (freckles):

so there were no exponents in the choices?

OpenStudy (kabase1234):

Yes all of the "n" on the end of each choice is an exponent

OpenStudy (freckles):

ok well if you didn't know how to answer any of the questions i presented above ... then try this: plug in 1 into all of your choices and see which one spots out the first term in your sequence (the -3)

OpenStudy (kabase1234):

How? I'm not good at these

OpenStudy (freckles):

how to replace n with 1 and evaluate?

OpenStudy (freckles):

wherever there is an n, you put 1

OpenStudy (freckles):

then see if you get the result a1=-3 from it

OpenStudy (freckles):

for example: you should see two n's here \[a_n=-4 \cdot 3^{n+1}\]

OpenStudy (freckles):

replace the n's with 1

OpenStudy (kabase1234):

Ohhhh okay

OpenStudy (kabase1234):

I got -3 as the first number am I right?

OpenStudy (kabase1234):

So it would either be B or C right?

OpenStudy (freckles):

which of those two choices given you -3 when you replace n with 1

OpenStudy (kabase1234):

B

OpenStudy (freckles):

sounds good you can remember it like this you have a common ratio that means you can always do term divided by previous term and get the same number that is you can do (term)/(previous term)=same number so you can always just plug your numbers into \[a_n=\text{ first term} \cdot (\text{common ratio number} )^{n-1}\]

OpenStudy (kabase1234):

Wow thank you so much so to clarify, I am correct with choice B?

OpenStudy (freckles):

that nth term right there only works for geometric sequences

OpenStudy (freckles):

"sounds good"<-----

OpenStudy (freckles):

well that is if you meant choice B to be a_n=-3 \cdot 4^{n-1}

OpenStudy (kabase1234):

Can you help with just one more? I can open a new question

OpenStudy (freckles):

\[a_n=-3 \cdot 4^{n-1}\]

OpenStudy (kabase1234):

Yes thank you so much